OpenStudy (anonymous):

Am I right? I think it's 108 or 72.... Sara is wrapping gift boxes. Each gift box is a rectangular prism with a square base of side 6 inches. The difference between the heights of the two gift boxes is 3 inches. If the larger box has a surface area of 216 square inches, what is the difference in the surface area of the two boxes?

OpenStudy (anonymous):

@009infinity

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (dmnts):

Height of the bigger box is a, then\[S _{b}=72+24a=216\]find a and insert it in the smaller boxs surface area formula: \[S _{s}=72+4(a-3)\]then do this: \[S _{b}-S _{s}\]

OpenStudy (dmnts):

Correction:\[S _{s}=72+24(a-3)\]

OpenStudy (dmnts):

Or you can just substract without finding a

OpenStudy (fibonaccichick666):

draw a picture

OpenStudy (dmnts):

The answer should be 72

OpenStudy (anonymous):

oh no wt i will draw it again

OpenStudy (fibonaccichick666):

@TheHero draw a pic then label said pic then write up the eq for surface area, then I will tell you if you are right :)

OpenStudy (anonymous):